OpenStudy (anonymous):

use implicit differentiation to find dy/dx of cos(x−y)+siny=√2

OpenStudy (anonymous):

dy/dx = 0

OpenStudy (anonymous):

are you sure? how did you get that?

OpenStudy (anonymous):

dy/dx=-f'y/f'x f'y already equals 0 so dy/dx=0

OpenStudy (anonymous):

differentiating with respect to x. -sin(x-y)(-dy/dx)+cosy dy/dx = 0 dy/dx (cosy + sin(x-y))=0 dy/dx = 0

OpenStudy (anonymous):

so what's the equation for the tangent line at \[(\pi/2, \pi/4)\]

OpenStudy (anonymous):

i guess that's what i got too I just didn't think it was right because now i don't get how to do the equation for the tangent line @u1umer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

hey could you help me find the equation of the tangent line?

OpenStudy (anonymous):

y= pie/4 equ of tangent line